OpenStudy (anonymous):

Use cross products to demonstrate whether or not the two ratios are equivalent:4/6 and 14/21

OpenStudy (anonymous):

How do i do this?

OpenStudy (anonymous):

if \[\frac{ 4 }{ 6} = \frac{ 14 }{ 21} \] then \[1= \frac{ 14*6 }{ 21*4} \]

OpenStudy (anonymous):

check to see if 14*6 = 21*4

OpenStudy (anonymous):

its equal :O

OpenStudy (anonymous):

TY so much!!!